# SparseLLM: Towards Global Pruning of Pre-trained Language Models

### v3

@@ Abstract
The transformative impact of large language models (LLMs) like LLaMA and GPT on natural language processing is countered by their prohibitive computational demands. Pruning has emerged as a pivotal compression strategy, introducing sparsity to enhance both memory and computational efficiency. Yet, traditional global pruning is impractical for LLMs due to scalability issues, while local pruning, despite its efficiency, leads to suboptimal solutions. Addressing these challenges, we propose SparseLLM, a novel framework that redefines the global pruning process into manageable, coordinated subproblems, allowing for resource-efficient optimization with global optimality. SparseLLM’s approach, which conceptualizes LLMs as a chain of modular functions and leverages auxiliary variables for problem decomposition, not only facilitates a pragmatic application on LLMs but also demonstrates significant performance improvements, particularly in high-sparsity regimes where it surpasses current state-of-the-art methods. Code can be found at https://github.com/BaiTheBest/SparseLLM.
@@ 1 Introduction
Large language models (LLMs) Touvron et al. (2023); OpenAI (2023) have recently transformed the field of natural language processing (NLP) by delivering exceptional results across a variety of intricate language benchmarks Wei et al. (2022); Bommarito II and Katz (2022); Bubeck et al. (2023). Nonetheless, these models, with billions of parameters, generally necessitate significant computational resources. To make LLMs more accessible, extensive efforts have been devoted to model compression of LLMs Xu and McAuley (2023); Bai et al. (2024), including pruning, quantization, knowledge distillation, and low-rank factorization. Pruning, by introducing sparsity, jointly enhances memory and computational efficiency and offers unparalleled flexibility, seamlessly integrating with any LLMs, thus standing out as a highly effective and widely adopted compression strategy.
Model pruning has a long history LeCun et al. (1989) and has proven effective in applications related to vision and smaller language models Hoefler et al. (2021). However, conventional pruning techniques, which rely on global pruning and require loading the entire model into the same GPU Mallya and Lazebnik (2018); Singh and Alistarh (2020), become impractical for today’s LLMs due to their vast size. Recently, several local pruning methods, which consider compressing each layer separately, and the overall compressed model is then obtained by “stitching together” the individually compressed layers, are proposed for billion-scale LLMs. SparseGPT Frantar and Alistarh (2023), an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieves up to 60% parameter reduction with minimal performance loss. Another approach, Wanda Sun et al. (2023), introduces a novel pruning criterion that evaluates weights by considering both magnitude and related input activations. Despite its efficiency gains, local pruning only aims to minimize the local error for each specific layer under sparsity constraints, resulting in a suboptimal solultion for the overall model. This is because local pruning over-aligns the intermediate layers’ activations, leading to suboptimal performance, especially in high-sparsity regimes Singh and Alistarh (2020); Sung et al. (2023).
@@ 2 Related work
Unstructured pruning looks at simplifying the complexity of LLMs by removing certain parameters regardless of the model’s inherent structure. This approach typically involves setting a threshold to nullify parameters below it, leading to a model with a non-uniform sparse structure. SparseGPT Frantar and Alistarh (2023), an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieves up to 60% parameter reduction with minimal performance loss. A novel pruning criterion is introduced in Wanda Sun et al. (2023), which evaluates weights by considering both magnitude and related input activations. This approach is beneficial in linear layers of LLMs, helping to identify and remove less significant weights. Tuli and Jha Tuli and Jha (2023) proposed DynaTran, a dynamic inference scheme for pruning activations at runtime, supported by a specially designed ASIC architecture, AccelTran, to enhance transformer inference throughput.
On the other hand, structured pruning involves the selective removal of groups of weights, where “group” might mean blocks of weights, filters, attention heads, or other structures conducive to hardware acceleration. Ma et al. Ma et al. (2023) introduced the LLM-Pruner, a framework designed for structured pruning of LLMs, which utilizes a combination of first-order data and Hessian information for effective importance estimation. This aids in identifying crucial groups for pruning. Li et al. Li et al. (2023) proposed LoSparse, a novel approach combining low-rank and sparse matrix approximations to balance pruning and expressive power. Tao et al. Tao et al. (2023) extended this concept to pruning hidden dimensions in LLMs, including embedding layers and attention heads. ZipLM Kurtic et al. (2023), a structured compression method for LLMs, is proposed to optimize for compression and accuracy while considering specific hardware constraints.
@@ 4.3 Algorithm design
Formally speaking, rather than try to solve Eq. 3 directly, we first relax the constraints by adding an $\ell_{2}$-penalty function to the objective and attack the unconstrained problem:
@@ 4.3.1 SparseLLM on OPT models
In this work, we consider the alternating method to optimize our Eq. 5, i.e., optimize each variable while keeping the rest fixed. The careful and elaborate design of our Eq. 5 allows us to derive a closed-form solution to every subproblem as shown in the following.
@@ 4.3.2 SparseLLM on LlaMA models
Consider the objective function given by Eq. 5, under the condition that the activation function $\phi$ is ReLU. Notice that (1) the objective function is convex with respect to each variable when all others are fixed, and (2) given that closed-form solutions exist for the subproblems in the alternating optimization scheme, the proposed algorithm is guaranteed to converge to a global optimum.
@@ 4.4 Time complexity analyses
The proposed SparseLLM consists of three main steps, with the overall time complexity being the sum of the complexities of these steps. In the weights pruning step, the complexity is dominated by the pseudo-inverse computation of matrix $\boldsymbol{a}_{\ell}$ (dimensions $n\times h$), which is $O(nh^{2})$. Using SparseGPT as the solver, the exact pruning step has a complexity of $O(h^{3})$. The second step, updating activation, involves matrix inversion of the weight matrix $\mathbf{W}_{\ell}$ (size $h\times h$) with a complexity of $O(h^{3})$. The third step, updating outputs, has a lower complexity. Thus, the overall algorithm complexity is bounded by $O(h^{3})$, making our method’s per-epoch time complexity comparable to SparseGPT.
@@ 5 Experiments
Experiments setup. We implement SparseLLM in PyTorch Paszke et al. (2019) and use the HuggingFace Transformers library Wolf et al. (2019) for handling models and datasets. All pruning experiments are conducted on NVIDIA A100 GPUs. For calibration data, we follow Frantar and Alistarh (2023) and use 128 2048-token segments, randomly chosen from the first shard of the C4 Raffel et al. (2020) dataset. This represents generic text data crawled from the internet and ensures our experiments remain zero-shot as no task-specific data is seen during pruning. We followed existing work Frantar and Alistarh (2023); Sun et al. (2023) and pruned all linear layers (in FFN and MHA) to the target sparsity.
Models, datasets & evaluation. We consider the OPT model family Zhang et al. (2022) and LlaMA-2 model family Touvron et al. (2023) in our experiments. We show results on different sizes of models to provide a broader picture for the performances of SparseLLM. In terms of metrics, we mainly focus on perplexity, which is known to be a challenging and stable metric that is well-suited for evaluating the accuracy of compression methods Yao et al. (2022); Dettmers and Zettlemoyer (2023). We consider the test sets of raw-WikiText2 Merity et al. (2016) (WT2) and PTB Marcus et al. (1994) as well as a subset of the C4 validation data, all popular benchmarks in LLM compression literature Yao et al. (2022); Park et al. (2022); Frantar and Alistarh (2023); Sun et al. (2023). For additional interpretability, we also provide zero-shot accuracy results following the same setup of Sun et al. (2023), which is based on the popular EleutherAI-eval harness Gao et al. (2023).
@@
5.1 Results and analyses
Pruning vs. model sizes. We begin by exploring the pruning capabilities of SparseLLM across various model sizes in comparison to baseline methods. For each model, we consider unstructured sparsity ranging from 70% to 90% with a 10% increment, as well as a 3:4 semi-structured sparsity. The 3:4 semi-structured sparsity is inspired by our preliminary results that suggest good performance SparseLLM at high sparsity regimes. However, note that two of our baselines, Magnitude and Wanda, are unable to be configured to this sparsity out-of-box. We conduct a sensitivity study on the calibration sample sizes (see Appendix A.3) and use calibration sample sizes between 32 and 64 for all experiments. Moreover, we prune the first 50% of the Transformer decoder layers in each model to achieve a balance between the computation resources and the performances. Detailed results can be found in Table 1 and Table 5 (see Appendix A.5). The perplexity results of the dense models are reported next to the names of the models.
From the table, it shows a general trend of increasing perplexity with increasing sparsity. Moreover, we observe a trend of decreasing perplexity for SparseGPT and SparseLLM at the same sparsity with increasing model sizes. However, such a trend is not obvious for Magnitude and Wanda. We also observe that SparseGPT and SparseLLM consistently outperform Magnitude and Wanda by a significant margin. For smaller sparsity, SparseLLM achieves comparable perplexity to SparseGPT. As we increase the sparsity, SparseLLM starts to demonstrate noticeable improvements over SparseGPT. In numerous instances for the OPT model family, SparseLLM achieves perplexity reductions of more than 50% compared to SparseGPT. We also see that performance improvements from SparseLLM over SparseGPT are more significant for the OPT model family than the LlaMA-2 model family.
Zero-shot experiments. To further conclude the evaluations and discussions, we show results for several zero-shot tasks in Table 2 and Table 6 (see Appendix A.5) comparing SparseGPT and SparseLLM. These evaluations are known to be relatively noisy Dettmers et al. (2022), but more interpretable. We also report the results for zero-shot tasks from the dense models in the “Dense" row. We see that the accuracy of both methods decreases with increasing sparsity, which is expected, as more parameters are pruned. A similar trend of increasing accuracy with increasing model size is observed too. Across all the tasks, OBQA and ARC-c remain the most challenging ones as the accuracy for both methods is 30% or below 30% while both methods perform well for BoolQ, RTE, WinoGrande, and ARC-e. In general, SparseLLM is able to achieve higher accuracy in the majority of tasks across the models of different sizes in both OPT and LlaMA-2 model families.
@@ 6 Conclusion
Our work presents SparseLLM, a cutting-edge framework poised to redefine the compression of LLMs through sparsity. By adeptly circumventing the scalability issues of global pruning and optimizing the local suboptimality of existing methods, SparseLLM stands as a significant advancement in the field. Our empirical results affirm its efficacy, particularly in high-sparsity environments, where it achieves a notable reduction in perplexity, thereby setting a new precedent for model compression. The versatility and minimal computational overhead of SparseLLM complement its integration with current pruning technologies, underscoring its potential as a universal tool for enhancing the performance and accessibility of LLMs.
@@ A.3 Calibration Samples
We study how the computation time of SparseGPT and SparseLLM varies with different model sizes, as illustrated in Table 3 and Table 4 for OPT models and LlaMA-2 models. The rate at which the time taken increases is comparable for SparseGPT and SparseLLM as the model size increases. Additionally, computation time for SparseLLM are reported for a configuration of 4 to 10 epochs. As we have reported in Section 5, SparseLLM can reduce the training loss in as few as 2 to 3 epochs. This suggests that the proposed SparseLLM remains computationally efficient.

### v4

@@ Abstract
The transformative impact of large language models (LLMs) like LLaMA and GPT on natural language processing is countered by their prohibitive computational demands. Pruning has emerged as a pivotal compression strategy, introducing sparsity to enhance both memory and computational efficiency. Yet, traditional global pruning is impractical for LLMs due to scalability issues, while local pruning, despite its efficiency, leads to suboptimal solutions. Addressing these challenges, we propose SparseLLM, a novel framework that redefines the global pruning process into manageable, coordinated subproblems, allowing for resource-efficient optimization with global optimality. SparseLLM’s approach, which conceptualizes LLMs as a chain of modular functions and leverages auxiliary variables for problem decomposition, not only facilitates a pragmatic application on LLMs but also demonstrates significant performance improvements, particularly in high-sparsity regimes, surpassing current state-of-the-art methods. Our source code is publicly available at https://github.com/BaiTheBest/SparseLLM.
@@ 1 Introduction
Large language models (LLMs) Touvron et al. (2023); OpenAI (2023) have recently transformed the field of natural language processing (NLP) by delivering exceptional results across a variety of intricate language benchmarks Wei et al. (2022); Bommarito II and Katz (2022); Bubeck et al. (2023). Nonetheless, these models, with billions of parameters, generally necessitate significant computational resources. To make LLMs more accessible, extensive efforts have been devoted to model compression of LLMs Xu and McAuley (2023); Bai et al. (2024a), including pruning, quantization, knowledge distillation, and low-rank factorization. Pruning, by introducing sparsity, jointly enhances memory and computational efficiency and offers unparalleled flexibility, seamlessly integrating with any LLMs, thus standing out as a highly effective and widely adopted compression strategy.
Model pruning has a long history LeCun et al. (1989) and has proven effective in applications related to vision and smaller language models Hoefler et al. (2021). However, conventional pruning techniques, which rely on global pruning and require loading the entire model into the same GPU Mallya and Lazebnik (2018); Singh and Alistarh (2020), become impractical for today’s LLMs due to their vast size. Recently, several local pruning methods have been proposed for billion-scale LLMs. These methods compress each layer separately, and the overall compressed model is then obtained by “stitching together” the individually compressed layers. SparseGPT Frantar and Alistarh (2023), an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieves parameter reduction of up to 60% with minimal performance loss. Another approach, Wanda Sun et al. (2023), introduces a novel pruning criterion that evaluates weights by considering both magnitude and related input activations. Despite its efficiency gains, local pruning only aims to minimize the local error for each specific layer under sparsity constraints, resulting in a suboptimal solution for the overall model. This is because local pruning over-aligns the intermediate layers’ activations, leading to suboptimal performance, especially in high-sparsity regimes Singh and Alistarh (2020); Sung et al. (2023).
@@ 2 Related work
Unstructured pruning Zhang et al. (2024); Bai et al. (2024b) looks at simplifying the complexity of LLMs by removing certain parameters regardless of the model’s inherent structure. This approach typically involves setting a threshold to nullify parameters below it, leading to a model with a non-uniform sparse structure. SparseGPT Frantar and Alistarh (2023), an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieves up to 60% parameter reduction with minimal performance loss. A novel pruning criterion is introduced in Wanda Sun et al. (2023), which evaluates weights by considering both magnitude and related input activations. This approach is beneficial in linear layers of LLMs, helping to identify and remove less significant weights. Tuli and Jha Tuli and Jha (2023) proposed DynaTran, a dynamic inference scheme for pruning activations at runtime, supported by a specially designed ASIC architecture, AccelTran, to enhance transformer inference throughput.
On the other hand, structured pruning involves the selective removal of groups of weights, where “group” might mean blocks of weights, filters, attention heads, or other structures conducive to hardware acceleration. Ma et al. Ma et al. (2023) introduced the LLM-Pruner, a framework designed for structured pruning of LLMs, which utilizes a combination of first-order data and Hessian information for effective importance estimation. This aids in identifying crucial groups for pruning. Li et al. Li et al. (2023) proposed LoSparse, a novel approach combining low-rank and sparse matrix approximations to balance pruning and expressive power. Tao et al. Tao et al. (2023) extended this concept to pruning hidden dimensions in LLMs, including embedding layers and attention heads. ZipLM Kurtic et al. (2023), a structured pruning method for LLMs, is proposed to optimize for compression and accuracy while considering specific hardware constraints. More recently, Xia et al introduced LLM-shearing Xia et al. (2023), a structured pruning method that scales down LLaMA models by selectively pruning layers, heads, and dimensions. This approach, combined with dynamic data batching, reduces pre-training compute costs while maintaining competitive performance, outperforming similar open-source models on key tasks.
@@ 4.3 Algorithm design
Formally speaking, rather than trying to solve Eq. 3 directly, we first relax the constraints by adding an $\ell_{2}$-penalty function to the objective and attack the unconstrained problem:
@@ 4.3.1 SparseLLM on OPT models
In this work, we consider the alternating method to optimize our Eq. 5, i.e., optimize each variable while keeping the rest fixed. The careful and elaborate design of our Eq. 5 allows us to derive a closed-form solution to every subproblem as shown below.
@@ 4.3.2 SparseLLM on LlaMA models
Consider the objective function given by Eq. 5, under the condition that the activation function $\phi$ is ReLU. Notice that (1) the objective function is convex with respect to each variable when all others are fixed, and (2) given that closed-form solutions exist for the subproblems in the alternating optimization scheme, the proposed algorithm resembles multiblock ADMM which has been shown to converge to in many applications.
@@ 4.4 Time complexity analyses
The proposed SparseLLM consists of three main steps, with the overall time complexity being the sum of the complexities of these steps. In the weights pruning step, the complexity is dominated by the pseudo-inverse computation of matrix $\boldsymbol{a}_{\ell}$ (dimensions $n\times h$), which is $O(nh^{2})$. Using SparseGPT as the solver, the exact pruning step has a complexity of $O(h^{3})$. The second step, updating activations, involves matrix inversion of the weight matrix $\mathbf{W}_{\ell}$ (size $h\times h$) with a complexity of $O(h^{3})$. The third step, updating outputs, has a lower complexity. Thus, the overall algorithm complexity is bounded by $O(h^{3})$, therefore making our method’s per-epoch time complexity comparable to SparseGPT.
@@ 5 Experiments
Experiments setup. We implemented SparseLLM in PyTorch Paszke et al. (2019) and use the HuggingFace Transformers library Wolf et al. (2019) for handling models and datasets. All pruning experiments are conducted on NVIDIA A100 GPUs. For calibration data, we follow Frantar and Alistarh (2023) and use 128 2048-token segments, randomly chosen from the first shard of the C4 Raffel et al. (2020) dataset. This represents generic text data crawled from the internet and ensures our experiments are zero-shot as no task-specific data is seen during pruning. We followed existing work Frantar and Alistarh (2023); Sun et al. (2023) and pruned all linear layers (in FFN and MHA) to the target sparsity.
Models, datasets & evaluation. We consider the OPT model family Zhang et al. (2022) and LlaMA-2 model family Touvron et al. (2023) in our experiments as well as the most recent LlaMA-3 model. We show results on different sizes of models to provide a broader picture for the performances of SparseLLM. In terms of metrics, we mainly focus on perplexity, which is known to be a challenging and stable metric that is well-suited for evaluating the accuracy of compression methods Yao et al. (2022); Dettmers and Zettlemoyer (2023). We consider the test sets of raw-WikiText2 Merity et al. (2016) (WT2) and PTB Marcus et al. (1994) as well as a subset of the C4 validation data, all popular benchmarks in LLM compression literature Yao et al. (2022); Park et al. (2022); Frantar and Alistarh (2023); Sun et al. (2023). For additional interpretability, we also provide zero-shot accuracy results following the same setup of Sun et al. (2023), which is based on the popular EleutherAI-eval harness Gao et al. (2023).
5.1 Results and analyses
Pruning vs. model sizes. We begin by exploring the pruning capabilities of SparseLLM across various model sizes in comparison to baseline methods. For each model, we consider unstructured sparsity ranging from 70% to 90% with a 10% increment, as well as a 3:4 semi-structured sparsity. The 3:4 semi-structured sparsity is inspired by our preliminary results that suggest good performance SparseLLM at high sparsity regimes. However, note that two of our baselines, Magnitude and Wanda, are unable to be configured to this sparsity out-of-box. We conduct a sensitivity study on the calibration sample sizes (see Appendix A.3) and use calibration sample sizes between 32 and 64 for all experiments. Moreover, we prune the first 50% of the Transformer decoder layers in each model to achieve a balance between the computation resources and the performances. Detailed results can be found in Table 1 and Table 2 as well as Table 8 in Appendix A.5. Note that in Table 2 for LlaMA-3 model, we only compare SparseGPT to the proposed SparseLLM. The perplexity results of the dense models are reported next to the names of the models.
Table 1: Perplexity of OPT models for sparsity $\geq 70\%$; the lower the perplexity, the better.
@@ 5.1 Results and analyses
Table 2: Perplexity of LlaMA models for sparsity $\geq 70\%$; the lower the perplexity, the better.
@@
LlaMA-3 8b (WikiText2 (WT2): 6.14; PTB: 11.18; C4: 9.45)
Sparsity 70% 80% 90% 3:4
Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
SparseGPT 22.37 36.56 30.53 72.87 113.95 79.86 214.68 261.18 198.34 96.75 107.52 102.11
SparseLLM 20.98 33.78 28.94 57.83 85.98 72.18 197.47 241.68 181.69 76.33 99.54 93.68
Table 3: Perplexity of 2:4 sparsity; the lower the perplexity, the better.
OPT-1.3b OPT-2.7b OPT-6.7b OPT-13b
Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
Magnitude 96.68 133.92 48.08 272.34 308.55 267.70 64.11 92.23 82.67 67.07 110.77 52.61
Wanda 15.63 24.04 18.23 13.66 21.67 16.10 11.86 18.54 14.77 10.33 15.35 12.54
SparseGPT 15.11 23.71 17.88 12.62 19.28 15.12 11.30 16.90 13.51 10.20 15.14 12.48
SparseLLM 14.97 23.40 17.67 12.62 19.28 15.12 11.07 16.73 13.42 10.20 15.14 12.41
From the tables, it shows a general trend of increasing perplexity with increasing sparsity. Moreover, we observe a trend of decreasing perplexity for SparseGPT and SparseLLM at the same sparsity with increasing model sizes. However, such a trend is not obvious for Magnitude and Wanda. We also observe that SparseGPT and SparseLLM consistently outperform Magnitude and Wanda by a significant margin. For smaller sparsity, SparseLLM achieves comparable perplexity to SparseGPT. As we increase the sparsity, SparseLLM starts to demonstrate noticeable improvements over SparseGPT. In numerous instances for the OPT model family, SparseLLM achieves perplexity reductions of more than 50% compared to SparseGPT. We also see that performance improvements from SparseLLM over SparseGPT are more significant for the OPT model family than the LlaMA-2 model family.
We provide additional set of perplexity results for a 2:4 semi-structured sparsity for a few OPT models in Table 3. We see that SparseLLM and SparseGPT generally outperform Magnitude and Wanda while SparseLLM has comparable if not slightly better performances compared to SparseGPT with the 2:4 semi-structured sparsity. Note that a 2:4 semi-structure sparsity is considered to be in low sparsity regime.
Zero-shot experiments. To further conclude the evaluations and discussions, we show results for several zero-shot tasks in Table 4 and Table 5 as well as Table 9 in Appendix A.5, comparing SparseGPT and SparseLLM. These evaluations are known to be relatively noisy Dettmers et al. (2022), but more interpretable. We also report the results for zero-shot tasks from the dense models in the “Dense" row. We see that the accuracy of both methods decreases with increasing sparsity, which is expected, as more parameters are pruned. A similar trend of increasing accuracy with increasing model size is observed too. Across all the tasks, OBQA and ARC-c remain the most challenging ones as the accuracy for both methods is 30% or below 30% while both methods perform well for BoolQ, RTE, WinoGrande, and ARC-e. In general, SparseLLM is able to achieve higher accuracy in the majority of tasks across the models of different sizes in both OPT and LlaMA-2 model families.
Table 4: Accuracy (%) of zero-shot tasks for OPT models; the higher the accuracy, the better.
OPT-13b
Sparsity Method BoolQ RTE HellaSwag WinoGrande ARC-e ARC-c OBQA Mean
Dense 65.87 57.76 52.44 66.02 67.82 33.46 28.62 53.14
70% SparseGPT 63.03 54.87 50.89 65.43 67.47 32.85 26.40 51.56
SparseLLM 63.85 55.23 50.73 65.67 66.46 31.83 27.20 51.57
80% SparseGPT 59.72 52.35 46.82 61.48 62.50 31.23 21.80 47.99
SparseLLM 60.89 53.07 46.19 62.12 62.21 30.38 23.00 48.27
90% SparseGPT 47.49 52.71 33.17 51.54 39.98 21.33 17.80 37.72
SparseLLM 53.43 52.71 38.19 52.96 46.68 25.26 17.40 40.95
3:4 SparseGPT 47.55 53.43 31.30 50.20 37.63 22.53 17.60 37.18
SparseLLM 51.13 52.35 38.51 55.96 49.24 24.83 21.40 41.92
@@
Table 5: Accuracy (%) of zero-shot tasks for LlaMA models; the higher the accuracy, the better.
@@ 6 Conclusion
Our work presents SparseLLM, a cutting-edge framework poised to redefine the compression of LLMs through sparsity. By adeptly circumventing the scalability issues of global pruning and optimizing the local suboptimality of existing methods, SparseLLM stands as a significant advancement in the field. Our empirical results affirm its efficacy, particularly in high-sparsity environments. It achieves a notable reduction in perplexity, thereby setting a new precedent for model compression. The versatility and minimal computational overhead of SparseLLM complement its integration with current pruning technologies, underscoring its potential as a universal tool for enhancing the performance and accessibility of LLMs.
Acknowledgments and Disclosure of Funding
This work was supported by the National Science Foundation (NSF) Grant No.1755850, No.1841520, No.1942594, No.2403312, No.2007716, No.2007976, No.1907805. This work was supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research, under Contract DE-AC02-06CH11357. This research used resources of the Argonne Leadership Computing Facility at Argonne National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under contract DE-AC02-06CH11357.
@@ A.3 Calibration Samples
We study how the computation time per layer of SparseGPT and SparseLLM varies with different model sizes, as illustrated in Table 6 and Table 7 for OPT models and LlaMA-2 models. The rate at which the time taken increases is comparable for SparseGPT and SparseLLM as the model size increases. Additionally, computation time for SparseLLM are reported for a configuration of 4 to 10 epochs. As we have reported in Section 5, SparseLLM can reduce the training loss in as few as 2 to 3 epochs. This suggests that the proposed SparseLLM remains computationally efficient.